OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 3x − y = −3 and passes through the point (4, −2).

OpenStudy (anonymous):

try to rewrite the given equation on the form y=mx+b, where m =slope. example: 2x-y=4 , we isolate the -y, so, 2x - 2x - y = 4 - 2x, will be: - y = 4 - 2x.

OpenStudy (anonymous):

then, divide both sides by - 1, y = 2x - 4. this is already in the form y=mx+b, so, m = 2.

OpenStudy (anonymous):

when you already have the slope, and the point, you can just use the y-y1=(slope)(x-x1), then, simplify.